### v2

# Uncertainty-Aware Measurement of Scenario Suite Representativeness for Autonomous Systems

Abstract
Assuring the trustworthiness and safety of AI systems, e.g., autonomous vehicles (AV), depends critically on the data-related safety properties, e.g., representativeness and completeness of the datasets used for their training and testing. Among these properties, this paper focuses on representativeness—the extent to which the scenario-based data used for training and testing, reflect the operational conditions that the system is designed to operate safely in, i.e., Operational Design Domain (ODD) or expected to encounter, i.e., Target Operational Domain (TOD). We propose a probabilistic method that quantifies representativeness by comparing the statistical distribution of features encoded by the scenario suites with the corresponding distribution of features representing the TOD, acknowledging that the true TOD distribution is unknown, as it can only be inferred from limited data. We apply an imprecise Bayesian method to handle limited data and uncertain priors. The imprecise Bayesian formulation produces interval-valued, uncertainty-aware estimates of representativeness, rather than a single value. We present numerical examples comparing the distributions of the scenario suite and the inferred TOD across operational categories (e.g., weather and road type) under dependencies and prior uncertainty. We estimate representativeness locally (between categories) and globally as an interval.
I Introduction
The rapid integration of Artificial Intelligence (AI) into autonomous systems has enabled machines to perform increasingly complex and safety-critical functions that were traditionally carried out by humans. In the context of autonomous vehicles and other intelligent systems, this transformation raises a pressing challenge: ensuring that the data-related safety properties governing AI-based decision-making components are sufficient to enable safe and reliable operation.
Table I summarises data-related safety properties defined in PD ISO/PAS 8800:2024 [9], DATA SAFETY (V3.6) [4], and ISO/IEC 5259-1 [2], which together specify key dataset qualities for AI and autonomous-system safety assurance.
TABLE I: Data-related safety properties from PD ISO/PAS 8800:2024 [9], DATA SAFETY (V3.6) [4], and ISO/IEC 5259-1 [2].
Source Data-Related Safety Properties
PD ISO/ PAS 8800 [9] Accuracy; Completeness; Correctness / Fidelity; Independence of datasets; Integrity; Representativeness; Temporality; Traceability; Verifiability.
DATA SAFETY [4] Integrity; Completeness; Accuracy; Timeliness / Temporality; Verifiability; Fidelity / Representation; Lifetime; Confidentiality; Availability.
ISO/IEC 5259-1 [2] Portability; Understandability; Auditability.
While properties such as accuracy, completeness, integrity, and traceability (Table I) concern local or procedural aspects of data quality, representativeness uniquely addresses the global, distributional relationship between a dataset and the real-world conditions it is intended to model. Thus, among those properties, representativeness—defined as “the distribution of data corresponds to the information in the environment of the phenomenon to be captured; it is free of biases” [9]—is of particular importance.
In existing literature and standards, the term environment carries related but distinct meanings. The notion of environment ranges from Operational Design Domain (ODD) distributions [11] to Target Operational Domain (TOD) specifications [8]. ISO 34503 [8] distinguishes between ODD, which defines the conditions under which an ADS is designed to operate safely, and TOD, which describes the real-world conditions the ADS may encounter and must handle safely. The standard allows different relationships between these domains, including partial overlap or one being broader than the other [^1] In practice, the TOD is often broader than the ODD.. We restrict representativeness assessment to the intersection $\mathrm{TOD}\cap\mathrm{ODD}$, considering only ODD-compliant scenarios. Representativeness is therefore evaluated by comparing the statistical distribution of features (i.e., scenario attributes such as road, weather, and traffic conditions) in the scenario suites with the corresponding distribution of features representing this restricted TOD [^2] For brevity, we will refer to them as the “scenario suite distribution” and the “TOD distribution”..
In the context of AI safety assurance, and following ISO/PAS 8800 [9], the term data encompasses all information used to train, test, and validate AI components. In the case of autonomous systems, this data is typically scenario-based, describing structured representations of the situations the system may encounter during operation. A study [10] interprets this concept operationally by treating structured scenarios as the fundamental data that may be used to train and test AI models (scenario suites). Each scenario captures a combination of operational variables such as weather, road type, illumination, etc. Assessing AI system safety involves evaluating data-related safety properties (Table I), which together determine the adequacy of the dataset. Among the key adequacy properties, this paper focuses on representativeness.
As mentioned earlier, representativeness can be formally expressed as a distributional discrepancy between the empirical scenario suite and the true but unknown TOD distribution. The TOD distribution is unknown because it represents the real-world operational conditions that the system will encounter, which can only be estimated from limited, potentially biased, or evolving observational data.
A key challenge lies in the multidimensionality of the TOD, which includes variables such as weather, road type, traffic density, and lighting, each spanning multiple levels (e.g., rainy/clear, urban/rural). Estimating a joint distribution from limited observations is difficult. If the occurrence frequencies of each category (e.g., clear–rural–light traffic–day, rainy–rural–light traffic–day, etc.) were known with certainty, one could deterministically fit a distribution to the data. In practice, however, such knowledge is rarely available, motivating a probabilistic approach that assigns distributions over category occurrences rather than fixed values (l.h.s of Fig. 1).
In contrast, estimating the empirical distribution of the scenario suites used is relatively straightforward, as these data are directly available. This corresponds to a descriptive statistical analysis of the observed dataset, rather than statistical inference about an unknown population (r.h.s of Fig. 1).
The central contribution of this work is therefore the estimation of the TOD joint distribution using a Bayesian framework (statistical inference, shown on the l.h.s of Fig. 1) well-suited for reasoning under data sparsity and uncertainty. Furthermore, we employ imprecise probability to represent uncertainty in the prior distribution used in the Bayesian inference. In general, our work formalizes representativeness as a distributional discrepancy between the empirical scenario suite and the “unknown” TOD distribution, and by deriving a Bayesian framework to quantify it together with its epistemic uncertainty (Fig. 1). Consequently, the resulting discrepancy metric is expressed as an uncertainty-aware estimate interval.
Figure 1: Schematic illustration of the representativeness evaluation process. The true population of TOD scenarios is unknown, and sampling from it introduces uncertainty, represented by imprecise probability envelopes around the estimated distribution. Representativeness is then assessed by comparing this uncertain TOD distribution with the known distribution of the scenario test suite.
There are different statistical metrics available for comparing probability distributions. In this work, we employ Total Variation Distance (TVD) and Jensen–Shannon Divergence (JSD) for this purpose.
The remainder of this paper is organized as follows. Section II reviews related work on dataset representativeness in AI safety assurance. Section III presents the proposed metric for quantifying representativeness. A numerical example illustrating its application is provided in Sec. IV. Section V discusses the results, and finally, Sec. VI concludes the paper.
II Related Work
This section reviews existing approaches to representativeness evaluation and imprecise probability to position the proposed framework.
II-A Comparison with Existing Methods on Representativeness
Recent work in automated driving has recognized that the representativeness of test scenarios is a key determinant of the credibility of safety evidence. Langner et al. [11] formalize representativeness as a statistical distance between two empirical distributions: the observed open-road test data and the target operational design domain (ODD) reference statistics. De Gelder et al. [5] apply this idea to the comparison between generated scenario data and real-world driving data. They use a combination of singular value decomposition (SVD) and kernel density estimation (KDE) to model the joint distribution of driving parameters, and evaluate how closely the generated scenarios approximate the empirical distribution of real-world data. Hiller et al. [7] shift focus from the metric itself to the process of representativeness assessment. Their framework compares distributions of recorded driving data against reference statistics derived from heterogeneous sources such as OpenStreetMap, meteorological services, traffic counters, and accident databases. In [10], representativeness is defined as the similarity between the empirical distribution of operational conditions in the scenario suite and the corresponding distribution in the TOD. Both are modeled as categorical probability distributions derived from ISO-standardized operational factors. At the assurance level, the JRC report on Trustworthy Autonomous Vehicles [6] identifies representativeness and data coverage as measurable properties required for trustworthy validation and testing. While this establishes representativeness as a normative requirement, it does not specify how to quantify it or account for uncertainty.
The above methods evaluate representativeness as a distance between distributions of two empirical datasets: generated versus real-world [5], test versus reference ODD data [11], or scenario suites versus TOD [10]. The real-world dataset is treated as a reliable surrogate for the true operational domain, assuming that sufficient data exist to capture its variability. In practice, however, this assumption may not hold when real world data are sparse or partially observed. Some studies, such as Gelder et al [5], highlight this issue and quantify how sampling bias in real-world testing affects the validity of validation results using frequency-based divergence measures. Our probabilistic framework (Bayesian inference) addresses this by comparing the empirical distribution of the scenario suite with an estimated approximation of the true, but unknown, TOD distribution. Instead of assuming that the TOD can be directly estimated from abundant real-world data, we treat its distribution as uncertain and propagate epistemic uncertainty into the representativeness metrics. Thus, our work extends deterministic distribution-matching approaches into a robust, uncertainty-aware framework.
II-B Imprecise Probability
While classical Bayesian methods provide a coherent mechanism for uncertainty propagation, they depend on the specification of a single prior distribution, which can be challenging to justify under sparse data. Imprecise probability frameworks [1] relax this assumption by representing uncertainty using set of distributions (credal sets) instead of a single one (Table II). Estimating such set of posterior distributions poses significant computational challenges. This is because there is generally no monotonic relationship between the prior and the posterior. However, when the prior belongs to the exponential family, a reparameterization can be applied—expressing the prior in terms of its strength and mean parameters—according to the iLUCK model proposed by [12] and applied in different studies ([13, 14]). In the context of our categorical data, the iLUCK model [12] allows the specification of bounds on these parameters. This formulation preserves conjugacy while enabling monotonic propagation of prior imprecision into posterior intervals. This monotonicity property allows us to compute the posterior bounds by evaluating only the corner cases of the prior parameter space, rather than calculating the posterior for every possible prior configuration.
TABLE II: Comparison of Classical Bayesian and Imprecise Probability for a coin-flip example.
Aspect Classical Bayesian Imprecise Probability
Prior $\mathrm{Beta}(3,3)$ $\mathrm{Beta}(\alpha,\beta),$ $\alpha,\beta\!\in\![2,4]$
Posterior $\mathrm{Beta}(8,10)$ $\mathrm{Beta}(5{+}\alpha,7{+}\beta)$, $\alpha,\beta\!\in\![2,4]$
Posterior mean $\mathbb{E}[\theta\mid D]=0.44$ [0.39, 0.48][0.39,\,0.48]
III A representativeness metric
Given the definition of TOD and tailored definition of data in this study, it is conceived as a multidimensional space that describes the range of real-world operating conditions under which an AI system is expected to function safely. Each dimension of this space corresponds to an operational variable, such as weather, road type, illumination, etc., (Table III).
Because many of these operational variables are continuous in nature—such as illumination intensity, temperature, or speed—it is often easier and more practical to represent the TOD in a finite, analyzable form through discretization (see Table III). In practice, standards and safety guidelines such as ISO/PAS 8800:2024 [9] and DATA SAFETY (V3.6) [4] establish the requirement for a well-defined coverage of the operational domain but leave the specific discretization boundaries to be determined by domain-specific standards or, in their absence, through expert judgment.
TABLE III: Illustrative hierarchy of variables and categories in TOD.
Hierarchy level Example elements Interpretation / role
Variables/ Factors Weather, Road type, Time of day, Traffic density, Speed band, etc. Dimensions describing operational and environmental conditions. Each variable may be continuous or discrete.
Discretized levels Weather: {Clear, Rain, Fog} Road type: {Highway, Urban, Rural} Time: {Day, Dusk, Night} Continuous variables are partitioned into discrete bins (e.g., speed ranges) to enable categorical representation.
Joint categories (Rainy, Urban, Night) (Clear, Highway, Day) (Foggy, Rural, Dusk), etc. Each unique combination of discretized variable levels forms one category $k\in\{1,\ldots,K\}$, representing a distinct operational condition in the TOD.
Given these joint categories of variables, our goal is to compare the estimated distributions of the TOD and the scenario suite to evaluate representativeness. In the first step, we sample from the TOD population and estimate its distribution using an imprecise Bayesian framework, as formulated in Theorem 1. Then in Sec. III-B a metric is presented to quantify the representativeness.
III-A Estimating the Posterior Distribution Envelopes of TOD
Theorem 1 formalizes the computation of lower and upper bounds on the TOD posterior under imprecise priors.
Theorem 1 (Posterior inference for the TOD distribution)
Let the Target Operational Domain (TOD) be represented by a categorical random variable $Z\in\{1,\ldots,K\}$ with unknown probability vector $\boldsymbol{\theta}_{\text{TOD}}=(\theta_{1},\ldots,\theta_{K})\in\Delta_{K-1}$ where, $\Delta_{K-1}$ is all valid probability distributions over $K$ categories:
$\Delta_{K-1}=\left\{\boldsymbol{\theta}=(\theta_{1},\ldots,\theta_{K})\in\mathbb{R}^{K}:\theta_{k}\geq 0,\;\sum_{k=1}^{K}\theta_{k}=1\right\}$
Given observed category counts $\mathbf{k}=(k_{1},\ldots,k_{K})$ from $n$ i.i.d.333i.i.d. means that each scenario is treated as an independent draw from the TOD distribution, so the outcome of one scenario does not affect another. samples, the likelihood is multinomial:
$f(\mathbf{k}\mid\boldsymbol{\theta}_{\text{TOD}})\propto\prod_{k=1}^{K}\theta_{k}^{k_{k}}.$
Assuming a Dirichlet prior with strength $n^{(0)}$ and prior mean vector $\boldsymbol{y}^{(0)}$, the posterior is
$\boldsymbol{\theta}_{\text{TOD}}\mid\mathbf{k}\sim\mathrm{Dirichlet}\!\big(n^{(0)}\boldsymbol{y}^{(0)}+\mathbf{k}\big).$
To account for prior uncertainty, let $\boldsymbol{y}^{(0)}\in\mathcal{Y}_{\text{TOD}}\subseteq\Delta_{K-1}$ (where, $\mathcal{Y}_{\text{TOD}}$ is the set of all possible prior mean vectors) and $n^{(0)}\in[\underline{n},\overline{n}]$. Then the posterior credal set is
$\mathcal{P}(\boldsymbol{\theta}_{\text{TOD}}\mid\mathbf{k})=$
$\hskip 14.22636pt\Big\{\mathrm{Dirichlet}\!\big(n^{(0)}\boldsymbol{y}^{(0)}+\mathbf{k}\big):\;\boldsymbol{y}^{(0)}\in\mathcal{Y}_{\text{TOD}},\;n^{(0)}\in[\underline{n},\overline{n}]\Big\}$
The joint posterior forms a family of Dirichlet distributions over the simplex, whose lower and upper envelopes are attained at the extreme points of $\mathcal{Y}_{\text{TOD}}$ and $\underline{n},\overline{n}]$. Marginally, each $\theta_{k}$ follows a corresponding family of Beta [^4] This is the mathematical property of the Dirichlet. distributions, with posterior bounds obtained in the same way.
See Appendix A for the outline of proof.
III-B Representativeness Metric under Uncertainty
Given the posterior TOD distribution from Theorem 1, representativeness is quantified as the discrepancy between $\hat{\boldsymbol{\pi}}$ (empirical scenario suite distribution) and $\boldsymbol{\theta}_{\mathrm{TOD}}\mid\mathbf{k}$. Representativeness is quantified by the alignment between these two categorical distributions over the same $K$ joint categories, using two complementary measures from information and probability theory:
III-B1 Total Variation Distance (TVD)
TVD provides an intuitive measure of the maximum difference in probability that the two distributions assign to any event:
$D_{\mathrm{TV}}(\hat{\boldsymbol{\pi}}_{\mathrm{suite}},\boldsymbol{\theta}_{\mathrm{TOD}})=\frac{1}{2}\sum_{k=1}^{K}|\hat{\pi}_{k}-\theta_{k}|$ (1)
The factor $\frac{1}{2}$ ensures that $D_{\mathrm{TV}}\in[0,1]$, where 0 indicates perfect alignment and 1 indicates complete disjointness. The TVD has a clear operational interpretation: it represents the maximum difference in probability that the two distributions assign to any subset of categories. A value of $D_{\mathrm{TV}}=0.05$, for example, means that the two distributions differ by at most 5% in their probability assignments to any collection of operational conditions.
III-B2 Jensen-Shannon Divergence (JSD)
JSD is a symmetrized and smoothed version of the Kullback-Leibler divergence, defined as:
$D_{\mathrm{JS}}(\hat{\boldsymbol{\pi}}_{\mathrm{suite}},\boldsymbol{\theta}_{\mathrm{TOD}})=$
$\hskip 14.22636pt\frac{1}{2}D_{\mathrm{KL}}(\hat{\boldsymbol{\pi}}_{\mathrm{suite}}\|{\bf m})+\frac{1}{2}D_{\mathrm{KL}}(\boldsymbol{\theta}_{\mathrm{TOD}}\|{\bf m}),$ (2)
where ${\bf m}=\frac{1}{2}(\hat{\boldsymbol{\pi}}_{\mathrm{suite}}+\boldsymbol{\theta}_{\mathrm{TOD}})$ is the average distribution, and $D_{\mathrm{KL}}$ denotes the Kullback-Leibler divergence:
$D_{\mathrm{KL}}({\bf p}\|{\bf q})=\sum_{k=1}^{K}p_{k}\log\frac{p_{k}}{q_{k}},\qquad p=\hat{\pi}_{\mathrm{suite}},\quad q=\boldsymbol{\theta}_{\mathrm{TOD}}$
III-B3 Propagating Epistemic Uncertainty
When using an imprecise prior, the posterior TOD distribution is represented by a credal set rather than a single distribution. Consequently, both $D_{\mathrm{TV}}$ and $D_{\mathrm{JS}}$ become interval-valued:
$D_{\mathrm{TV}}$ $\in\left[\underline{D}_{\mathrm{TV}},\overline{D}_{\mathrm{TV}}\right]\qquad D_{\mathrm{JS}}\in\left[\underline{D}_{\mathrm{JS}},\overline{D}_{\mathrm{JS}}\right]$
where the lower and upper bounds are obtained by evaluating the discrepancy measure at the boundary values of the prior strength parameter $n^{(0)}\in[\underline{n},\overline{n}]$. Because the posterior mean $\boldsymbol{\theta}_{\mathrm{TOD}}$ varies monotonically with the prior strength $n^{(0)}$, and both $D_{\mathrm{TV}}$ and $D_{\mathrm{JS}}$ are continuous in $\boldsymbol{\theta}_{\mathrm{TOD}}$, the minimum and maximum discrepancy values over the prior interval $\underline{n},\overline{n}]$ occur at the boundary points of this interval.
These bounds illustrate epistemic uncertainty in representativeness, indicating how sensitive the assessment is to prior beliefs about the TOD. Wide intervals suggest the need for more operational data to refine the estimate.
IV Numerical Example
The numerical example presented in this section is intentionally simplified and uses a small number of abstract categorical variables to illustrate the proposed method in a transparent and reproducible manner.
IV-A Problem Setting: Common Cases
The TOD and the scenario suite share the same variable set, discretization, and categorical space. Both are defined over five operational factors: $W,R,D,T,S,$ where, Weather $W\in\{\text{Clear},\text{Adverse}\}$, Road $R\in\{\text{Highway},\text{Urban}\}$, Time $D\in\{\text{Day},\text{Night}\}$, Traffic $T\in\{\text{Light},\text{Heavy}\}$, and Speed $S\in\{\leq 50,\;>50\}$. Each combination $(w,r,d,t,s)$ defines one of $K=2^{5}=32$ categories (Table IV), which serve as a common comparison basis for both the TOD and the scenario suite.
TABLE IV: Mapping of joint categories to 5-bit codes.
Code Weather Road ToD Traffic Speed
0 Clear Highway Day Light $\leq 50$
1 Clear Highway Day Light >50>50
2 Clear Highway Day Heavy $\leq 50$
3 Clear Highway Day Heavy >50>50
$\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$
31 Adverse Urban Night Heavy >50>50
To ensure realism and consistency, dependencies among these factors are explicitly incorporated in both models through a shared Bayesian-network structure $(R\!\to\!W,\;D\!\to\!T,\;R\!\to\!S)$. This structure captures operationally relevant relationships such as weather conditions depending on road type, traffic patterns varying with time of day, and speed distributions differing across road environments. This structure is applied consistently to both the scenario suites and TOD. It captures realistic relationships among scenario variables: road type influences both weather exposure and typical speed, while time of day affects traffic conditions. The root marginal distributions are specified as
$P(R)=[0.60,\,0.40]\quad(\text{Highway, Urban}),$
$P(D)=[0.65,\,0.35]\quad(\text{Day, Night})$
The corresponding conditional probability tables (CPTs) are summarized in Table V.
TABLE V: Conditional Probability Tables (dependent TOD prior).
$P(W\mid R)$   (Weather $\mid$ Road type)
Clear Adverse
Highway 0.75 0.25
Urban 0.625 0.375
$P(T\mid D)$   (Traffic $\mid$ Time of day)
Light Heavy
Day 0.65 0.35
Night 0.50 0.50
$P(S\mid R)$   (Speed $\mid$ Road type)
$\leq 50$ >50>50
Highway 0.30 0.70
Urban 0.65 0.35
IV-B Problem Setting: Scenario Suite
The scenario suite serves as an empirical test distribution for comparison with the inferred TOD distribution, emulating autonomous-vehicle scenario generation where rare but safety-critical conditions are intentionally oversampled to ensure sufficient evaluation coverage.
A synthetic dataset of n=1,600n=1{,}600 scenarios was generated by incorporating dependencies among operational factors $(R\!\to\!W,\;D\!\to\!T,\;R\!\to\!S)$. To reflect operational risk emphasis, higher sampling weights were assigned to more challenging conditions — adverse weather ($\times 1.4$), urban roads ($\times 1.3$), nighttime ($\times 1.25$), heavy traffic ($\times 1.2$), and low speeds ($\leq 50~\mathrm{km/h},\times 1.15$). Within each conditional relationship, multipliers were applied within parent contexts and then renormalized to preserve valid conditional probability tables. For example, in the conditional distribution $P(W\mid R)$, the weight for adverse weather was increased by $1.4$ relative to clear conditions, and the resulting probabilities were normalized within each road type.
The joint probability of each category $c=(w,r,d,t,s)$ in the scenario suite was computed as
$\hat{\pi}_{\text{suite}}(c)\;\propto\;P(R{=}r)\,\times P(D{=}d)\times$
$\hskip 8.5359pt\,P(W{=}w\mid R{=}r)\times\,P(T{=}t\times\mid D{=}d)\times\,P(S{=}s\mid R{=}r)$
where each conditional term reflects both the dependency structure and applied weighting multipliers. After normalization, the probabilities form a valid joint distribution:
$\hat{\pi}_{\text{suite}}=\frac{\hat{\pi}_{\text{suite}}(c)}{\sum_{c}\hat{\pi}_{\text{suite}}(c)}.$
Integer scenario counts were then obtained as
$k(c)=\operatorname{round}\bigl(1600\cdot\hat{\pi}_{\text{suite}}(c)\bigr),\quad\text{with}\quad\sum_{c}k(c)=1600.$
The resulting $\hat{\pi}_{\text{suite}}$ thus constitutes a dependency-aware empirical distribution, reflecting both realistic correlations among operational factors and deliberate oversampling of safety-critical conditions. By design, this distribution differs from the TOD prior $\boldsymbol{\theta}_{\mathrm{TOD}}$, introducing a controlled mismatch that enables the representativeness analysis to quantify how these oversampling strategies affect alignment with the expected real-world operational domain.
As shown in Fig. 1, the numerical example illustrates three distributions: the empirical scenario-suite distribution obtained directly from the 1600 generated scenarios; the unknown true TOD distribution representing real operational exposure; and the imprecise TOD distribution, given by the interval-valued posterior inferred from limited operational samples and prior assumptions. Representativeness is assessed by comparing the scenario-suite distribution with this uncertain TOD distribution.
IV-C Problem Setting: TOD
The prior mean vector $\boldsymbol{y}^{(0)}$ over the 32 joint categories was defined assuming dependencies among the five operational factors, represented through a Bayesian network with three directed links: $R\rightarrow W,\quad D\rightarrow T,\quad R\rightarrow S.$ Given this dependency structure, the joint prior mean is computed as
$y^{(0)}_{(w,r,d,t,s)}=P(R{=}r)\,\times P(D{=}d)\,\times$
$\hskip 8.5359ptP(W{=}w\mid R{=}r)\,\times P(T{=}t\mid D{=}d)\,\times P(S{=}s\mid R{=}r)$
Normalization ensures that $\sum_{w,r,t,d,s}y^{(0)}_{(w,r,t,d,s)}=1$.
The prior strength is modeled as imprecise,
$n^{(0)}\in[\underline{n},\overline{n}]=[5,\,20],$
representing uncertain confidence in the baseline distribution.
Posterior inference
Given observed category counts $\mathbf{k}=(k_{1},\ldots,k_{32})$, the posterior distribution is
$\boldsymbol{\theta}_{\mathrm{TOD}}\mid\mathbf{k},n^{(0)}\sim\mathrm{Dirichlet}\!\big(n^{(0)}\mathbf{y}^{(0)}+\mathbf{k}\big),$
with posterior mean per category (for illustration purpose):
$\mathbb{E}[\theta_{k}\mid\mathbf{k},n^{(0)}]=\frac{n^{(0)}y_{k}^{(0)}+k_{k}}{n^{(0)}+n}.$
Evaluating this expression at $n^{(0)}=\underline{n}$ and $n^{(0)}=\overline{n}$ yields the lower and upper posterior expectations, defining credible bounds on the TOD distribution.
IV-D Numerical results
Local representativeness analysis across joint operational categories
Fig. 2 shows two probability distributions defined over the same 32 categories:
Figure 2: Comparison between the empirical scenario suite distribution and the posterior mean of the TOD distribution (assuming $n^{(0)}=10$).
Each pair of points represents one unique combination of operational conditions (e.g., weather, road type, time of day, traffic, and speed). Although this is not a joint comparison aggregated across variables, it provides a fine-grained, category-level view of how closely the scenario suite aligns with the inferred TOD distribution.
The near overlap between the orange (suite) and blue (posterior TOD) points indicates a strong alignment for most categories. Small deviations, typically for adverse or nighttime conditions, reflect deliberate over-sampling in the scenario suite design. This visualization highlights which operational categories contribute most to representativeness discrepancies before summarization with global metrics (e.g., TVD, JSD).
A high discrepancy indicates that the scenario suite does not reflect real operational conditions. Per-category differences can identify under- and over-represented situations, guide scenario generation or reduction, and iteratively refine the suite until the discrepancy falls below an acceptable level.
Joint-distribution representativeness
Beyond individual categories, the global similarity between the scenario-suite distribution $\hat{\boldsymbol{\pi}}_{\mathrm{suite}}$ and the expected posterior TOD distribution [^5] Depending on the comparison objective, it is also possible to compare the lower-bound (conservative) and upper-bound distributions. $\mathbb{E}\!\left[\boldsymbol{\theta}_{\mathrm{TOD}}\mid\mathbf{k},\,n^{(0)}\right]$ is quantified using the TVD (1) and the JSD (III-B2). These measures capture the overall mismatch between the two 32-dimensional joint distributions, independently of the category ordering. The values $D_{\mathrm{TV}}=0.00227$ and $D_{\mathrm{JS}}=0.00001$ (at $n^{(0)}=10$) indicate that the scenario suite is highly representative of the inferred TOD, with negligible probability redistribution across conditions.
Over the prior-strength interval $n^{(0)}\in[5,\,20]$, the distances vary within $D_{\mathrm{TV}}\in[0.00114,\,0.00452]$ and $D_{\mathrm{JS}}\in[0.00000,\,0.00004]$. The TVD shows an overall mismatch below 0.5% (Maximum distance $\times 100=0.00452\times 100\simeq 0.5$). The JSD further shows that the two distributions carry almost identical information content: they demonstrate that the scenario suite remains globally representative of the operational conditions characterized by TOD.
Effect of prior values
The results in Tables  VI and VII quantify how the TVD between the scenario suite and the posterior TOD distribution varies with the prior strength parameter $n^{(0)}$ in the imprecise model.
TABLE VI: TVD for different fixed prior-strength values.
$n^{(0)}$ 5 10 20
$D_{\mathrm{TV}}$ 0.0011 0.0023 0.0045
As the prior strength $n^{(0)}$ increases, the posterior distribution becomes more influenced by the prior and less by the observed scenario frequencies, leading to a gradual increase in $D_{\mathrm{TV}}$ VI. This monotonicity caused by iLUCK model [12] reflects how epistemic confidence in the prior shifts the posterior away from the empirical scenario suite (small width of the TVD intervals, e.g., $[0.0031,0.0045]$, for small width of $n^{(0)}$, e.g., $n^{(0)}\in[15,20]$).
While the numerical example considers a scenario suite broadly aligned with the inferred TOD, the proposed metrics also capture strong mismatch: when scenarios are generated independently of the TOD structure or uniformly at random, TVD and JSD increase toward their upper bounds, indicating poor representativeness and enabling identification of highly non-representative suites.
TABLE VII: Effect of the length of the interval of $n^{(0)}\in[a,b]$.
$\underline{n}^{(0)}$ $\overline{n}^{(0)}$ $D_{\mathrm{TV}}(\underline{n}^{(0)})$ $D_{\mathrm{TV}}(\overline{n}^{(0)})$ $\underline{D}_{\mathrm{TV}}$ $\overline{D}_{\mathrm{TV}}$
5 20 0.0011 0.0045 0.0011 0.0045
10 20 0.0023 0.0045 0.0023 0.0045
15 20 0.0031 0.0045 0.0031 0.0045
V Discussion
V-A Limitations of single-valued representativeness metrics
As discussed in Section II, existing approaches assume the reference distribution is known with sufficient confidence. In data-sparse settings, this is problematic: single-valued representativeness metrics conceal sensitivity to prior assumptions, whereas interval-valued results reveal the range of plausible conclusions. These are not meant to replace single-valued metrics, but to complement them when data are limited or uncertain by exposing epistemic uncertainty that single-point results would otherwise hide.
V-B Prior distribution
In applying the proposed model, the main task is to specify the prior mean vector $\boldsymbol{y}^{(0)}$ and the prior strength $n^{(0)}$, which together represent the analyst’s belief about the TOD before observing data. Each element of $\boldsymbol{y}^{(0)}$ represents the expected frequency of a TOD condition (for example, how often “rainy–urban–night” situations occur), while $n^{(0)}$ quantifies the analyst’s confidence in this belief.
In practice, the prior mean $\boldsymbol{y}^{(0)}$ is not fixed but inferred from various forms of background knowledge. When sufficient operational or simulation data are available, the empirical proportions of observed conditions can be used directly as an estimate of $\boldsymbol{y}^{(0)}$. In more data-sparse contexts, expert elicitation provides an alternative: domain specialists may indicate how often specific conditions are likely to occur (for example, “most operations occur on highways during daytime”). These qualitative statements are translated into numerical probabilities that sum to one. In cases where experts can only specify approximate ranges rather than precise probabilities, these ranges define upper and lower bounds for the corresponding prior mean components.
In addition, published reference statistics—such as national traffic surveys, meteorological archives, or the multi-source environmental datasets used in Hiller et al. [7]—can provide baseline frequency estimates for relevant factors like weather type, road category, or traffic density. When no credible information exists, a uniform prior $\boldsymbol{y}^{(0)}=(1/K,\ldots,1/K)$ expresses a state of complete ignorance by assigning equal probability to all possible categories.
V-C Dependence and Independence among TOD Categories
In this study, the TOD is defined as a combination of discrete operational factors such as weather, road type, illumination, and traffic density etc. To derive the joint distribution, it is necessary to account for the dependencies that may exist among these factors, as such factors are rarely independent in real-world operational environments. Certain weather conditions may be more frequent on rural roads than in urban areas; night-time driving is typically associated with lower traffic density; and fog or heavy rain may be confined to specific geographical regions or seasons. These dependencies imply that the true TOD distribution $\boldsymbol{\theta}_{\mathrm{TOD}}$ cannot, in general, be expressed as a simple product of marginal probabilities:
$P(\text{Road},\text{Weather},\text{Time})\neq P(\text{Road})\,P(\text{Weather})\,P(\text{Time}).$
Neglecting such correlations can lead to systematic under- or over-representation of certain operational categories in the scenario suite, biasing the representativeness assessment.
In Sec. IV, dependencies were explicitly modeled to reflect realistic operational conditions. These dependencies need to be determined based on prior expert knowledge and supported by available historical data, which indicate systematic relationships between certain environmental and contextual factors. If the dependencies are not well understood and independence between variables must be assumed, a sensitivity analysis can help to avoid misleading conclusions by quantifying the potential impact of this assumption on the representativeness metrics.
VI Summary and Conclusions
This paper presented a probabilistic framework for assessing the representativeness of scenario datasets with respect to the TOD. Unlike existing approaches based on fixed distributions and single-point discrepancy measures, the proposed approach models the TOD as an uncertain categorical distribution over joint operational conditions and applies a robust Bayesian formulation with imprecise Dirichlet priors to capture epistemic uncertainty about both prior beliefs and data sufficiency. The resulting posterior credal set enables transparent quantification of uncertainty in representativeness metrics (e.g., Total Variation Distance) between the observed scenario suite distribution and the inferred TOD distribution.
Unlike the numerical example, real-world ODDs involve many variables; the framework therefore relies on structured or factorized representations capturing only relevant dependencies.
Future work will extend the current framework along several directions. First, we will relax the i.i.d. assumption to account for potential correlations between scenarios, which may arise from temporal, spatial, or operational dependencies in real-world driving data. Next, we will address the dependence between the TOD and scenario suite distributions, acknowledging that both often stem from the same limited real-world dataset. Treating them as fully independent can conceal shared biases and limit the ability to reveal representational gaps. We also aim to relax the simplifying assumptions made in this study by incorporating continuous operational factors, thereby moving beyond purely discretized representations toward a hybrid framework that combines discrete and continuous variables. Beyond representativeness, we will develop quantitative metrics for other data-related safety properties—such as completeness, consistency, and relevance—and explore their interrelations within a unified framework. Ultimately, these assessments should be embedded within an Assurance 2.0 case (e.g., [3]), linking probabilistic evidence to structured, auditable safety arguments for trustworthy AI. While the numerical example demonstrates the feasibility and interpretability of the approach, broader empirical comparisons with existing representativeness metrics on larger and more realistic datasets are left for future work.
Acknowledgment
This work is supported by the European Union’s EU Framework Program for Research and Innovation Europe Horizon (grant agreement No 101202457) and Innovate UK through the DriveSafeAI Project (10062594). XZ’s contribution is supported by the UK EPSRC New Investigator Award [EP/Z536568/1]. SK’s contribution is supported by the UKRI Future Leaders Fellowship Grant [MR/S035176/1]. Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or European Research Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Proof Sketch of Theorem 1
Model setup
Let $Z\in\{1,\dots,K\}$ denote the TOD category with probability vector $\boldsymbol{\theta}=(\theta_{1},\ldots,\theta_{K})\in\Delta_{K-1}$. Given $n$ i.i.d. observations with count vector $\mathbf{k}=(k_{1},\ldots,k_{K})$, the likelihood is multinomial:
We assume that the TOD variable $Z$ takes one of $K$ mutually exclusive categories, with probabilities $\boldsymbol{\theta}=(\theta_{1},\ldots,\theta_{K})$, where $\sum_{k=1}^{K}\theta_{k}=1$. If we observe $n$ independent samples $z_{1},\ldots,z_{n}$ from this categorical distribution and record the number of occurrences in each category as $\mathbf{k}=(k_{1},\ldots,k_{K})$, then the joint probability of the observed counts follows the multinomial distribution:
[…]
Since the factorial term does not depend on $\boldsymbol{\theta}$, it can be dropped when forming the likelihood function for inference:
$L(\boldsymbol{\theta}\mid\mathbf{k})\propto\prod_{k=1}^{K}\theta_{k}^{\,k_{k}}.$
Thus, the multinomial likelihood correctly represents the sampling process for counts over multiple mutually exclusive categories. The sufficient statistic for $\boldsymbol{\theta}$ is indeed the count vector $\mathbf{k}$, since (by the factorization theorem) all information in the sample relevant to $\boldsymbol{\theta}$ is contained in these category counts.
Dirichlet prior and iLUCK reparameterization.
Assume a Dirichlet prior $\boldsymbol{\theta}\sim\mathrm{Dirichlet}(\boldsymbol{\alpha})$ with parameters $\boldsymbol{\alpha}=n^{(0)}\boldsymbol{y}^{(0)}$, where $\boldsymbol{y}^{(0)}$ is the prior mean and $n^{(0)}>0$ is the prior strength (equivalent sample size). The Dirichlet distribution is the conjugate prior of the multinomial likelihood. That means if the likelihood of the data given the parameters is multinomial, then the posterior distribution over the parameters has the same functional form (Dirichlet), which allows closed-form Bayesian updating. This is the iLUCK reparameterization, satisfying $\mathbb{E}[\theta_{k}]=y^{(0)}_{k}$ and $\alpha_{k}=n^{(0)}y^{(0)}_{k}$.
Precise Bayesian update
By conjugacy of the Dirichlet–multinomial pair,
$p(\boldsymbol{\theta}\mid\mathbf{k})\;\propto\;p(\mathbf{k}\mid\boldsymbol{\theta})\,p(\boldsymbol{\theta})\;\propto\;\prod_{k=1}^{K}\theta_{k}^{\,(\alpha_{k}^{(0)}+k_{k}-1)}.$
Hence, $\boldsymbol{\theta}_{\text{TOD}}\mid\mathbf{k}\;\sim\;\mathrm{Dirichlet}\!\big(n^{(0)}\boldsymbol{y}^{(0)}+\mathbf{k}\big),$ and each component marginally follows: $\theta_{k}\!\mid\!\mathbf{k}\sim\mathrm{Beta}(a_{k},a_{0}-a_{k})$ with $a_{k}=n^{(0)}y^{(0)}_{k}+k_{k}$ and $a_{0}=n^{(0)}+n$.
Imprecise Bayesian posterior.
Represent prior uncertainty by the set $\boldsymbol{y}^{(0)}\in\mathcal{Y}_{\text{TOD}}\subseteq\Delta_{K-1}$ and $n^{(0)}\in[\underline{n},\overline{n}]$. By conjugacy, each prior in this set yields a corresponding Dirichlet posterior, forming the posterior credal set:
$\mathcal{P}(\boldsymbol{\theta}_{\text{TOD}}\mid\mathbf{k})=$
$\hskip 14.22636pt\big\{\mathrm{Dirichlet}\!\big(n^{(0)}\boldsymbol{y}^{(0)}+\mathbf{k}\big):\,\boldsymbol{y}^{(0)}\in\mathcal{Y}_{\text{TOD}},\,n^{(0)}\in[\underline{n},\overline{n}]\big\}$
Because the Dirichlet update is affine in $(\boldsymbol{y}^{(0)},n^{(0)})$, the posterior set remains convex.
Posterior envelopes and extreme points
$\Delta_{K-1}$. Under the iLUCK [12] monotonicity result, the lower and upper bounds of this joint posterior family (the posterior envelope) are attained at the extreme points of $\mathcal{Y}_{\text{TOD}}\times[\underline{n},\overline{n}]$. Monotonicity holds because, under the iLUCK reparameterization, the Dirichlet shape parameters
$a_{k}=n^{(0)}y_{k}^{(0)}+k_{k},\qquad a_{0}=n^{(0)}+n,$
depend affinely on the prior mean and strength. This affine relationship preserves convexity of the credal set under updating, ensuring that posterior expectations, densities, and CDFs vary monotonically with $y^{(0)}$ and $n^{(0)}$. Consequently, the lower and upper Dirichlet posterior densities— that is, the joint posterior envelope over $\Delta_{K-1}$— are obtained by evaluating the Dirichlet density at the extreme prior configurations. Marginally, each component $\theta_{k}$ follows a Beta distribution whose CDF is the regularized incomplete Beta function $I_{t}(a_{k},a_{0}-a_{k})$, so the same boundary evaluation yields tight marginal bounds on posterior beliefs. The Beta marginal is correct because each component of a Dirichlet distribution has a univariate Beta distribution when integrating out the remaining dimensions, preserving exact consistency between the joint and marginal posterior forms. Hence, the marginal posterior envelopes can be estimated in the same way, by evaluating the Beta distributions at the same extreme prior configurations.